For the year ended December 31, 20Y8, ABC Inc. made total sales of $1,000,000, but expects to receive requests for refunds of returned or damaged goods that are 2% of total sales. The company also expects to receive returns of goods costing $12,000. Prepare the adjusting entries required at year-end if 20Y8 is the company’s first year of operations.
Answer #1

Damaged goods = 1000000 * 2%

= $20,000

Expected returns = $12000

Total sales returns for year end = 20000 + 12000 = $32000

Journal entry:-

Sales returns $32,000
Accounts receivables $32,000
( To record sales returns)